Circular No. 9096

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 W2 (BOATTINI) A. Boattini reports his discovery of a comet (discovery observation

tabulated below) on CCD images taken with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey; he notes a moderately condensed coma ∼ 7′′ wide, with no tail visible. R. E. Hill reports that four co-added 60-s CCD images taken in 2′′ seeing with the Mt. Lemmon 1.5-m reflector on Nov. 23.5 UT show a bright nuclear condensation with a small (1′′–2′′) coma and a short, broad tail 6′′–7′′ long in p.a. ∼ 80^{o}; his similar exposures on Nov. 24.44–24.46 show a 4′′ × 5′′ nuclear condensation that is elongated east-west with a 1′′ coma surrounding, and a broad, diffuse 30′′ tail in p.a. 90^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, W. H. Ryan writes that CCD images taken by E. V. Ryan and himself with the Magdalena Ridge Observatory 2.4-m f/8.9 reflector on Nov. 24.5 show a tail in p.a. ∼ 90^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Nov. 23.47597 6^{h}21^{m}30.^{s}50 +39^{o}58′41′′.9 19.5 Boattini

The available astrometry (including observations by Boattini on Nov. 21), the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2009-W103.

T = 2010 Mar. 23.937 TT  ω = 118.^{o}066
Ω = 199.323 }2000.0
q = 6.95031 AU  i = 164.512

COMET 169P/NEAT B. G. Marsden, Smithsonian Astrophysical Observatory, has identified

a comet found in STEREO spacecraft images with comet 169P. K. Battams, Naval Research Laboratory, reports that the comet appears very clearly (marginally diffuse with no tail) at apparent mag ∼ 9–10 on Nov. 12.9 UT in HI-1B images (limiting mag ∼ 13.5), when it was only ∼ 4^{o}.5 from the sun (r = 0.69 AU, Δ = 0.88 AU, and 0.38 AU from STEREO), so that forward scattering is a likely explanation for 169P’s being well above its observed brightness from ground-based observations (cf. MPEC 2009-W102). Over the next few days, 169P moved to increasing elongations, with its brightness falling gradually; by Nov. 21 (elongation 18^{o}), STEREO images show the comet to be noticeably smaller and fainter (mag perhaps 11–12).

2009 November 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green